# InstructPipe: Building Visual Programming Pipelines with Human Instructions

### v1

@@ 4.1. Pipeline Representation
The Visual Blocks system represents a pipeline as a Directed Acyclic Graph (DAG) in JSON format [^1] Pipeline in JSON files: see supplementary materials for examples.. To compress the verbose JSON file, InstructPipe represents pipelines as pseudocode (Gupta and Kembhavi, 2023; Surís et al., 2023), which can be further compiled into a JSON-formatted pipeline. The pseudocode representation is highly token-efficient. Figure 5 shows an example in which the pseudo compresses the JSON-based pipeline representation (2.8k tokens) into a 123-token pseudocode representation. The efficiency does come with a cost: it loses some fine-grained annotations of each node like property values (e.g., layout of the nodes, parameters of a segmentation model, and the degree of blurring parameters in an “image processor” node). InstructPipe leaves the task of finetuning these parameters to the user and focuses on generating the graphic structure of a visual programming pipeline.
Figure 5 provides an example of a pipeline (4(a)) and its corresponding pseudocode (4(b)). The syntax design is inspired by TypeScript. The structure is inspired by how academic papers present pseudocode (Zhang et al., 2021): an algorithm block typically starts with specifying the input/output and then explains the intermediate processor. We highlight the first line under the processor module (i.e., the operation of the PaLI node) in 4(b) in four different colors, representing four different components in the programming language. “pali_1_out” represents the output variable name of the node. “pali_1” is the unique ID of the node. The green symbol after the colon, i.e., “pali”, specifies the type of the node. In this example, the node with the ID of “$pali\_1$” is a “pali” node. The rest part in the bracket, i.e., “image=input_image_1, prompt=input_text_1”, defines the arguments of the nodes. In the input pseudocode, we do not annotate the output variable (i.e., there are no red colors in the highlighted line under the input module) because all the input nodes only export one value, and the output variable name is automatically annotated as the same symbol as the node id (i.e., “input_text_1”). Note that InstructPipe generates the text input (i.e., the property value) in the “input text” node. Therefore, the argument in “text=“caption this image in detail”” does not indicate that the “input_text” node accepts input edges, but accepts the node property input.
[Uncaptioned image] Figure 6. The prompt structure for the Node Selection module. Each node description is formated as ”{node types}: {short descriptions of the nodes}; {recommended node(s)}”. The node recommendation is optional.
[Uncaptioned image] Figure 7. The prompt structure for the Code Writer module. Detailed node configurations, see supplementary materials for examples, are listed in the highlighted region.
@@ 4.2. Node Selector
Node Selector aims to filter unrelated nodes by providing an LLM with a short description of each node. Figure 6 shows the prompt we use in Node Selector. Followed by a general task description and guidelines, we list all the node types with a short description that explains the function of the node. Several nodes come with recommendation(s) when the users interact with Visual Blocks, and we also include such node recommendations information in the prompt. The main intuition of this prompt design is based on how the existing open-source libraries (e.g., Numpy (Harris et al., 2020)) present a high-level overview of all functions [^2] See an example in the following link: https://numpy.org/doc/1.25/reference/routines.array-manipulation.html. The documentation typically presents a list of supported functions (in each category), followed by a short description so that developers can quickly find their desired functions. At the end of the prompt, we provide a list of Q&A as few-shot examples to support the LLM to learn and adapt to the context of the task.
@@ 4.3. Code Writer
With a pool of selected nodes, the Code Writer module can write pseudocode for rendering a target pipeline. Figure 7 shows the structure of the prompt utilized in this LLM module. Similar to section 4.2, the prompt starts with a general introduction and several guidelines. The major difference in the prompt design in this stage lies in the granularity of each node introduction. We provide a detailed configuration for each selected node with additional information, including 1) input data types, 2) output data types, and 3) an example, represented in pseudocodes, showing how this node connects to other nodes. We put a detailed explanation of the full node configuration in the supplementary materials. Similar to the previous LLM module (section 4.2), the prompt design here is also inspired by the documentation of several popular code libraries. Specifically, we gain inspiration from low-level function-specific documentation [^3] See an example in numpy.shape: https://numpy.org/doc/1.25/reference/generated/numpy.shape.html#numpy-shape, which typically includes a detailed description and data types in the input/output, followed by one or more examples of how developers can use this function with a few lines of codes.
The prompt also includes a list of Q&A as few-shot examples. However, providing few-shot examples in this stage is non-trivial. The reason lies in the dynamics of the node selection pool: a combination of all the nodes causes many possible selections, and it is impossible to design a list of few-shot examples for each possibility. Therefore, we only created an example list for each pipeline tag (i.e., “language”, “visual”, and “multimodal”) and intended to utilize these few-shot examples to cover most of the use cases. The intuition behind this design is: the in-context examples serve to list possible creative ideas of the pipelines in each tag, which can be condensed into few-shot examples rather than traversing all the possible combinations. This implies that the in-context examples may include nodes that are not selected in the prompts. According to our preliminary tests, such out-of-scope nodes cause negative effects on the generation results: LLMs tend to also “invent” new nodes that do not exist in our system, causing failure in node rendering. Note that we also observed this issue when combining the prompts used in (Gupta and Kembhavi, 2023). We eliminate this issue by using the prompt contents between the node configurations and the in-context examples (i.e., the contents start with “the following is a full list of …” in Figure 7). This aims to inform LLMs of the exceptions for invention.
((a)) Before layout optimization.
((b)) After layout optimization.
Figure 8. A comparison of the same generated pipeline before and after layout optimization. (a): Each node is assigned with a default property value when convert pseudocode into a JSON file for rendering the pipeline. Such default values will cause sub-optimal visual effects. (b): Our layout optimization process re-arranges the layout for better presentation of the pipeline.
@@ 4.4. Code Interpreter
Finally, InstructPipe employs a code interpreter to parse the generated pseudocode and compile a JSON-formatted pipeline with an automatic layout. We delineate the graph compilation and rendering procedure below:
@@
    Lexical Analysis: InstructPipe first tokenizes each line of the pseudocode is into output variable id, node id, node type, and node arguments (section 4.1).
@@
    Graph Generation: The tokenized results allow us to build a graph structure in a node-graph editor that connects each node as specified by the pseudocode. We then generate JSON-formatted code, which Visual Blocks uses to render the pipeline in the node-graph editor. Note that the JSON code includes far more parameters than those used in InstructPipe for defining the graphic structure. InstructPipe fills texts in the “Input Text” node based on LLM-generated pseudocode and uses default values for the rest nodes. For example, by default, the temperature and the max output tokens for the PaLM node are set to 0.5 and 256, respectively. If users are not satisfied with the default values, they can interactively adjust the parameters in the node-graph editor.
@@
    Graph Rendering and Optimization: A problem with employing the default parameters is that it results in a chaotic distribution of nodes in the node-graph editor: each node is located in a predefined position, and the edges go across the workspace without a reasonable arrangement. Therefore, InstructPipe traverses the graph with a breadth-first search (BFS) and arranges the nodes based on their depth values in the DAG. As shown in Figure 8, such post-processing can effectively enhance the visualization of a generated pipeline.
Our exploration shows that LLMs may not always generate accurate code as expected. One typical issue is that Code Writer tends to invent illegal nodes out of our library of 27 nodes. To address this, InstructPipe disposes such lines that leverage illegal nodes. If legal nodes use the output of such an illegal node, InstructPipe discards the edge connection between them to prevent bugs when running the pipeline. Another common issue is that the generated code may not have a correct order. For example, the generated code may be “Line 1; Line 3; Line 4; Line 2;”, in which “Line 1 – 4” represents four lines of code, and the correct order should be from “Line 1” to “Line 4”. Such ordering issues would cause an input value of the current line to become undefined, because its definition is mistakenly placed afterward. We address this issue by stacking the line that is not ready for execution, similar to graph construction by traversing an adjacency list. The Code Interpreter runs in a loop to interpret the code until all legal lines in the stack become ready.
@@ 5. Technical Evaluation
Evaluating our system is challenging. Ideally, the evaluation set should cover 1) diverse visual programming pipelines (i.e., the pipeline factor) and 2) a variety of instructions created by different individuals for each pipeline (i.e., the human factor). However, such a combination requires a significant number of user evaluation sessions, which is impractical to deploy. For example, if an evaluation set includes 50 pipelines and 20 instructions for each pipeline from 20 different participants, it requires 1,000 user study sessions. Asking participants to attend multiple pipeline creation sessions seems efficient, but it introduces new learning transfer effects to the study. The learn effects usually require additional user study design with counterbalancing, which eventually increases the total number of required sessions.
To address this issue, we decouple the two factors that cause variations by decomposing the whole evaluation into two evaluation sessions. The major benefit of this design is that it allows us to deploy the evaluation tasks with a reasonable workload while maintaining the rigorousness of the whole evaluation. In our first evaluation session (i.e., technical evaluation), we focus on understanding the technical performance of InstructPipe among a variety of pipelines. In the second evaluation session (i.e., user evaluation), we control the pipeline factor and examine the human factor by recruiting different participants for prototyping controlled pipelines. The first technical evaluation provides crucial data on the “accuracy” of InstructPipe, which guides us to select such representative pipelines with as little bias as possible. More importantly, the user evaluation further allows us to understand how user experiences are affected under two conditions: with and without InstructPipe.
In this section, we focus on explaining the first step of our evaluation, the technical evaluation, and the following user evaluation is covered in the next section.
@@ 5.1. Data Collection
To conduct the technical evaluation, we first organized a two-day hybrid workshop with 23 participants, aiming to collect pipelines that Visual Blocks users build for their creative usage. Six attendees claimed that they had prior experience in using Visual Blocks. At the beginning of the workshop, we gave a 15-minute tutorial walking the participants through the nodes and the pipeline-building process using Visual Blocks. After the tutorial, attendees created pipelines independently. We required the participants to export the JSON file from Visual Blocks and upload it to our Google Drive every time they finished a pipeline. We also emphasized that they should caption their pipelines when they export the files. Such data pairs (pipeline/caption) by the participants constitute important raw data we use for the technical evaluation.
@@
Note that Visual Blocks includes more primitive nodes than the 27 nodes covered by InstructPipe. The workshop is an open-ended creation process, in which participants are free to any node available in Visual Blocks. The InstructPipe feature was not available in the workshop.
5.2. Data Post-Processing
Figure 9. Example pipelines participants built in the workshops.
  • Isolated graphs. There exist pipelines that include at least one isolated subgraph. The isolated subgraph, as opposed to the main graph, is defined as a graph (or a node) that has no connection to the main graph in the pipeline that provides the main functionality of the pipeline. The “Image viewer” node on the bottom-left corner of 7(b) We observed that some participants typically would like to explore the system by working on a separate sub-space. While we acknowledge its usefulness, leaving such “redundant” graphs in the raw data for the evaluation would cause issues when we calculate the number of user interactions (i.e., the metric used in the evaluation that will be defined in the next subsection).
The observation motivated us to post-process the raw data to present more rigorous evaluation results. We first removed incomplete pipelines and the isolated graphs in each pipeline (if there are any). Although the low-quality captions may reflect real-world scenarios on how the users may use our system, such a hypothesis cannot be officially validated. In our data collection process, participants are situated in the scenarios of writing captions instead of instructing an AI assistant. When users generate instruct pipelines, they may be aware of the importance of writing good instructions, and thus, such low-quality captions would not be as usual. More importantly, the technical evaluation focuses on considering the variation of the pipeline, instead of text instructions. In the next user evaluation (section 6), we will explicitly situate a new group of participants in instructing the assistant and observe the user behaviors and the variation of human instructions on the same pipelines. Therefore, we were motivated to enhance the quality of the captions for fair technical evaluations. Two authors individually annotated the caption of each pipeline separately by referring to the original captions and pipelines authored by the participants. It is important to note that we finished the workshop and the data annotation task before we completed the system implementation. The two authors had no experience using InstructPipe before completing the annotation. We believed this process could effectively enhance the quality of the captions while maintain fairness of the technical evaluation.
As we clarified in section 5.1, the workshop is designed to be an open-ended creation process. This indicates that the dataset inevitably includes out-of-scope nodes like “custom scripts” (in which the participants write code to process the input data and return custom outputs; see 8(b) for an example) and “TFLite model runner” (which call a custom TensorFlow model with a URL input of the model in the TF-Hub). As we explained in section 4.1, InstructPipe focuses on generating the graphic structure of a pipeline without considering the property value within each node. Different from the majority of nodes in Visual Blocks, nodes like “custom scripts” and “TFLite model runner” derive their intrinsic functionality from the property values. Without this information, the entire pipeline’s rationale is obscured. Therefore, we removed pipelines that contained “out-of-scope” nodes in our technical evaluation.
The final 48 pipelines (out of 64 pipelines) are comprised of 23 language pipelines, seven visual pipelines, and 18 multi-modal pipelines. Figure 9 shows three pipelines created by the participants. 8(b) is an example of the pipelines that include out-of-scope nodes, and therefore is not included in the final 48 pipelines. In the technical evaluation, we ran our generation algorithm on the pipeline captions six times (three times for each caption $\times$ two captions from two authors for each pipelines) and evaluated the generation results using the metric that will be introduced below.
@@ 5.3. Metric: The Number of User Interactions
Evaluating a generated pipeline is more complex than other tasks that have a universally recognized definition of accuracy. In this project, intuitively, an accurate generation implies that users only need to do very few edits (or even no edits) based on the generated result. This intuition inspires our definition of “the number of user interactions” as the core metric used in our evaluations:
@@
Note that there are countless ways to modify a generated pipeline toward a complete pipeline in practice. Nevertheless, the minimal number of user interactions, representing the “smartest” way(s) to make a modification, is deterministic, and this is an objective metric that can fairly reveal how many minimal efforts users need to spend to achieve their goal. A pipeline is considered complete when it satisfies the given instruction. We calculate the number of interactions from the sum of two events: 1) adding or deleting a node and 2) adding or deleting an edge between nodes. Note that when a node with edges connected is deleted, our system will auto-remove these edges. In such instances, we only register one interaction for the node deletion.
Our definition of the number of user interactions has two important implications. First, a complete pipeline after user interaction does not need to be the same as the corresponding pipeline in the dataset. As long as it fulfills the task described in the caption, we consider the pipeline complete. Second, our definition does not consider interactions of modifying property values of a node, e.g., typing in a text box or selecting a value in a drop-down box. We argue that such interactions are highly node-dependent and are hard to quantify objectively. More importantly, as we explain in section 4.1, the generation of property values is out of the scope of this work.
In the technical evaluation with various pipelines, it is unfair to report an averaged absolute value of user interactions because the complexity of the pipelines varies dramatically. For instance, the user may need three edits based on a generated result to complete a large pipeline that requires 20 edits from scratch. In another pipeline, the user also needs to do three edits starting from the generated result, but the whole pipeline only takes three edits to finish. Averaging these absolute values does not provide reasonable insights into how accurate the generation is. Therefore, we reported an averaged ratio of user interactions required to complete a pipeline “from our generated pipeline” to that “from scratch” as our target metric in the technical evaluation.
@@
5.4. Results
Table 1 summarizes the results of the technical evaluation. Compared to building a pipeline from scratch, InstructPipe allows the user to finish a pipeline with 18.9% of the user interactions (as defined in section 5.3), demonstrating the effectiveness of the InstructPipe support. Seven generated pipelines directly satisfied with instructions without user interactions in all six trials, and 38 generated pipelines completed at least once in any of the six trials.
@@ 6. User Evaluation
While the technical evaluation demonstrates the accuracy of InstructPipe among various real pipelines created by participants, it is still unclear what is the actual user experience when real users go through the entire system workflow. Additionally, there is a lack of subject variation on a controlled group of pipelines in the technical evaluation. Therefore, we conducted an in-person user study of InstructPipe with another group of participants, aiming to provide more insights into our system performance. The study recruitment was in accordance with the ethics board of our institution. We obtained participant consent before the study began.

### v2

@@ 4.1. Pipeline Representation
The Visual Blocks system takes a JSON-format data as input and renders a directed acyclic graph (DAG) in the visual programming workspace. Therefore, the ultimate goal of InstructPipe is to generate the JSON file; however, directly generating the long JSON file is computationally expensive. For example, the JSON file for rendering the pipeline in 4(a) contains approximately 2.8k tokens. To address this issue, we utilize the pseudocode representation of a DAG, and define this token-efficient data format as the output data format of our LLM module. 4(b) shows the corresponding pseudocode representation of the pipeline in 4(a), and the it only contains 123 tokens. The pseudocode representation simply stores the DAG information of a visual programming pipeline without other information such as node parameters (e.g., the “max_tokens” configuration of an LLM module) and the layouts of the nodes. This indicates that InstructPipe leaves the task of node parameter tuning to the user, which we believe is a more novice-friendly task, and focuses on providing technical assistants on selecting and connecting nodes.
In the following content, we provide detailed explanation on the pseudocode design and implementation. As we mentioned above, Fig. 4 provides an example of a pipeline (4(a)) and its corresponding pseudocode (4(b)). The syntax design is inspired by TypeScript, and the overall structure is inspired by how academic papers present pseudocode (Zhang et al., 2021) in an algorithm block. In 4(b), we highlight the first line under the processor module (i.e., the operation of the PaLI node) in different colors, representing four different components in the programming language. “pali_1” is the unique node ID. The green symbol after the colon, i.e., “pali”, specifies the node type. In this example, node ID “$pali\_1$” is a “pali” node. The arguments in brackets, i.e., “image=input_image_1, prompt=input_text_1”, specifies the input variables (or input edges in the graph) of this node. “pali_1_out” represents the output variable name. For input nodes, the output variable name is the same as the node id, so we do not annotate the output variable with a separate name (e.g., “input_image_1: input_image()” instead of “input_image_1 = input_image_1: input_image()”). Note that InstructPipe generates texts (i.e., the node parameter) in the “input text” node. Therefore, the argument in “text=“caption this image in detail”” does not indicate that the “input_text” node accepts input edges, but accepts the node parameter input as a special case.
Figure 5. The prompt structure for the Node Selection module. Each node description is formated as ”{node types}: {short descriptions of the nodes}; {recommended node(s)}”. The node recommendation is optional.
Figure 6. The prompt structure for the Code Writer module. Detailed node configurations, see the appendix for examples, are listed in the highlighted region.
@@ 4.2. Node Selector
Node Selector filters out unrelated nodes by providing the LLM with a short description of each node. Fig. 5 shows the prompt we use in Node Selector. Followed by a general task description and guidelines, we list all node types with a short description that explains the function of each. Several nodes come with recommendation(s) when the users interact with Visual Blocks, and we also include such node recommendations in the prompt. The main intuition of this prompt design is based on how existing open-source libraries (e.g., Numpy (Harris et al., 2020)) present a high-level overview of all functions [^2] See an example in the following link: https://numpy.org/doc/1.25/reference/routines.array-manipulation.html. The documentation typically presents a list of supported functions (in each category), followed by a short description so that developers can quickly find their desired functions. At the end of the prompt, we provide a list of Q&As as few-shot examples to support the LLM to learn and adapt to the context of the task.
@@ 4.3. Code Writer
With a pool of selected nodes, the Code Writer module can write pipeline rendering pseudocode. Fig. 6 shows the structure of the prompt utilized in this LLM module. Similar to section 4.2, the prompt starts with a general introduction and several guidelines. The major difference in the prompt design in this stage lies in the granularity of each node introduction. We provide a detailed configuration for each selected node with additional information, including 1) input data types, 2) output data types, and 3) an example, represented in pseudocode, for how this node connects to other nodes. We include a detailed explanation of the full node configuration in Section B.1.2. Similar to the previous LLM module (section 4.2), the prompt design here is also inspired by the documentation of existing software libraries. Specifically, we gain inspiration from low-level function-specific documentation [^3] See an example in numpy.shape: https://numpy.org/doc/1.25/reference/generated/numpy.shape.html#numpy-shape, which typically includes 1) a detailed description, 2) data types in the input/output, followed by 3) one or more examples of a few lines of code for how developers can use the function.
The prompt also includes a Q&A list as few-shot examples. However, providing few-shot examples in this stage is non-trivial. The reason lies in the dynamics of the node selection pool. A combination of all the nodes causes many possible options, and it is impossible to design a dedicated list of few-shot examples in each possible case. Therefore, we only created an example list for each pipeline tag (i.e., “language”, “visual”, and “multimodal”) and intended to utilize these few-shot pipelines to teach LLMs example use cases in each category. This implies that in-context pipelines may include nodes that were not selected for the prompt. This can potentially lead to LLM hallucinations (Huang et al., 2023b), i.e., utilizing the nodes that do not exist in our node library. We mitigated this issue by adding specific prompts that explicitly show a list of supported nodes (i.e., the contents start with “the following is a full list of …” in Fig. 6). However, LLM hallucination is a community-wise challenge, and we also find that our approach cannot completely eliminate this issue in visual programming. Therefore, InstructPipe conduct a sanity check for the Code Writer outputs and directly disposes of the line of pseudocode with such hallucinated nodes. This can ensure that the generated code is in a valid data format for rendering the pipeline in Visual Blocks.
@@ 4.4. Code Interpreter
After our LLM modules generate the pseudocode, InstructPipe employs a code interpreter to parse the generated pseudocode and compile a JSON-formatted pipeline with an automatic layout. Since we incorporated standard approaches to achieving such conversion from the pseudocode to the JSON file, which we do not intend to claim as our main contributions, we briefly summarize our implementation into the following three steps for simplicity and elaborate low-level implementation details at Appendix B.2:
@@
    Lexical Analysis: InstructPipe first tokenizes each line of the pseudocode into output variable id, node id, node type, and node arguments (section 4.1).
@@
    Graph Generation with Default Node Parameters: We generated a DAG based on the tokenized results, and apply predefined default node parameters in each generated node. For example, by default, the temperature and the max output tokens for the PaLM node are set to 0.5 and 256, respectively. If users are not satisfied with the default values, they can interactively adjust the parameters in the node-graph editor.
@@
    Layout Optimation: When pseudocode is converted into a JSON file, default node parameters will cause sub-optimal visual effects (11(a)). InstructPipe conducts a layout optimization process using the breadth-first search (BFS) algorithm, which re-arranges the layout for better presentation of the pipeline (11(b)).
@@ 5. Technical Evaluation
InstructPipe contributes a framework for generating specifications for visual programming pipelines based on text prompts from users. To characterize the system’s performance, we designed a technical evaluation to asses the accuracy of the generated graphs for a variety of prompts.
@@ 5.1. Data Collection
To compute the accuracy of our generated pipelines, we need to collect a corpus with pairs of instructions and their corresponding ground-truth pipelines. Therefore, we organized a two-day hybrid workshop with 23 participants, aiming to collect real pipelines that Visual Blocks users would build for their creative usage. The 23 participants (F: 6; M: 17) are composed of five software engineers, four research scientists, four students, three designers, two project managers, and two engineering mangers. Six attendees claimed that they had prior experience in using Visual Blocks. As this was a data collection study rather than a user study, where each participant here served as a data creator and annotator, we did not restrict participation to individuals who self-identified as novices. The workshop began with a 15-minute Visual Blocks tutorial walking the participants through the nodes and the pipeline-building process. After the tutorial, attendees created pipelines independently. Once they finished creating the pipelines, participants were required to caption their pipelines and upload them. We utilized this corpus of data pairs (caption/pipeline) as the data set for the technical evaluation.
@@
The workshop was an open-ended creation process, in which participants were free to use any node available in Visual Blocks with more than the 27 nodes covered by InstructPipe. The InstructPipe feature was not available in the workshop. After the workshop, we post-processed our collected data and achieved 48 pipelines (23 language pipelines, seven visual pipelines and 18 multi-modal pipelines) for our technical evaluations. The post-processing procedure details are presented in Appendix C.1.
@@ 5.2. Metric: The Number of User Interactions
To quantify the efficacy of InstructPipe based on our goal of accelerating and streamlining pipeline creation, we defined the metric Number of User Interactions as follow:
@@
This definition is mainly inspired by Graph Edit Distance (GED) in graph theory (Gao et al., 2010). Note that there are countless ways to modify a generated pipeline toward a complete pipeline in practice. Nevertheless, the minimal number of user interactions is deterministic, and this is an objective metric that can fairly estimate the amount of effort users need to spend to achieve their goal. A pipeline is considered complete when it satisfies the given instruction. We calculate the number of interactions across two types of events: 1) adding/deleting a node, and 2) adding/deleting an edge between nodes. In the technical evaluation, we report the average ratio of user interactions required to complete a pipeline “from our generated pipeline” compared to “from scratch” as our target metric. For example, if it takes 3 interactions to complete a pipeline from our generated results and takes 10 interactions to complete from scratch, then the ratio of interactions is $30\%$. Appendix C.2 contains further discussion of this metric.
@@
5.3. Experiment Setups and Results
We ran our generation algorithm on the pipeline captions six times (three times for each caption $\times$ two captions for each pipeline), and computed an averaged performance among the six trials for each pipeline. Table 1 summarizes the results of the technical evaluation. Compared to building a pipeline from scratch, InstructPipe allows the user to complete a pipeline with 18.9% of the user interactions, demonstrating the potential of InstructPipe to require more than 5X fewer interactions. Seven generated pipelines directly satisfied with instructions without user interactions in all six trials, and 38 generated pipelines completed at least once in any of the six trials.
@@ 6. User Evaluation
While the technical evaluation demonstrates the accuracy of InstructPipe among various real pipelines created by participants, it is still unclear what is the actual user experience when real users go through the entire system workflow. Therefore, we conducted an in-person user study of InstructPipe with another group of participants, aiming to provide more insights into our system performance as well as explore new user experiences brought by InstructPipe. The study recruitment was in accordance with the ethics board of our institution. We obtained participant consent before the study began.